User Tools

Site Tools


Recording to ATNF V013B.

Problem with X-band Tsys.

Observing stopped at 08:30UT to allow investigation. JER and SWA investigated and concluded that it looked like a cabling fault with no waveform getting to the RX to cycle the noise diode. Stowed and Ettore went to focus cabin to recable by taking the connection from the 13mm RX (which is not needed for this run) and connect it to Mars. Tsys is once again working and observing resumed at 09:35:20UT.

A fringe test at 14:43:50UT gave fringes on all baselines, however, the Parkes polarisations are crossed. I am almost 100% certain this happened during the Tsys investigation so the data prior to 08:30UT should be labelled correctly, where as the data since 09:35:20UT will incorrect. Need to swap at correlation.

lbaops/lbamay2010/v253bpalog.txt · Last modified: 2015/12/18 16:38 by